# Population Transformer: Learning Population-level Representations of Intracranial Activity

### v1

@@ Abstract
We present a self-supervised framework that learns population-level codes for intracranial neural recordings at scale, unlocking the benefits of representation learning for a key neuroscience recording modality. The Population Transformer (PopT) lowers the amount of data required for decoding experiments, while increasing accuracy, even on never-before-seen subjects and tasks. We address two key challenges in developing PopT: sparse electrode distribution and varying electrode location across patients. PopT stacks on top of pretrained representations and enhances downstream tasks by enabling learned aggregation of multiple spatially-sparse data channels. Beyond decoding, we interpret the pretrained PopT and fine-tuned models to show how it can be used to provide neuroscience insights learned from massive amounts of data. We release a pretrained PopT to enable off-the-shelf improvements in multi-channel intracranial data decoding and interpretability, and code is available at https://github.com/czlwang/PopulationTransformer.
1California Institute of Technology
{gchau, sabera, ssoedarm, yyue}@caltech.edu
2MIT CSAIL, CBMM
{czw, vsub851, boris, abarbu}@mit.edu
@@ 1 Introduction
Building effective representations of neural recordings is an important tool in enabling neuroscience research. We are particularly interested in modeling intracranial recordings, which rely on probes placed within the brain to provide high temporal resolution recordings of local neural activity [1, 2]. Because of its dispersed placement within the brain volume, intracranial recordings suffer from data sparsity. Moreover, there is often significant variability in probe placement across subjects [1, 2], leading to high variability in input channel meaning. Historically, constructing decoders from intracranial data has relied on supervised learning [3, 2, 4, 5, 6], but this requires experimenters to collect annotated data, which is scarce due to patient availability and labor-intensive labeling.
To improve decoding data-efficiency, self-supervised pretraining on unannotated data can be employed to first learn generic representations of the recordings. This means that the model does not have to use valuable annotated samples to learn how to do feature extraction before it can do classification, improving the reach of neuroscientific research.
@@
In this paper, we are interested in developing generic representations of multi-channel intracranial recordings that enable efficient adaptation to a wide range of downstream decoding tasks. Prior work has shown how to pretrain subject-specific [7] or channel-specific [8] models of intracranial data, but such techniques ignore inter-channel relationships or commonalities that might exist across subjects. The most general approach would be to pretrain using data from multiple datasets, but would require tackling the aforementioned challenges of sparse electrode coverage and variable electrode placement between subjects.
@@
We propose Population Transformer (PopT), a self-supervised pretraining approach on transformers [9] that learns subject-generic representations of arbitrary electrode ensembles (Figure 1). During pretraining, we simultaneously optimize both a channel-level and ensemble-level objective, that requires the model to (1) build representations of channels in the context of other channels and (2) meaningfully distinguish temporal relationships between different ensembles of channels.
@@
Our approach builds on top of existing single-channel embeddings, such as BrainBERT [8], which has two key advantages. First, by separating the single-channel embedding and multi-channel-aggregation into different modules, we make our approach agnostic to the specific type of temporal embedding used, leaving room for future independent improvements along either the temporal or spatial dimension, an approach that has been validated in video modeling [10]. Second, by taking advantage of the temporal feature-extraction learned during single-channel pretraining, we make our population-level training more data-efficient.
@@
Empirically, we find that our pretrained PopT outperforms non-pretrained aggregation approaches, highlighting the usefulness of learning spatial relationships during pretraining. Moreover, we find that these benefits hold even for subjects not seen during pretraining, lending to its usefulenss for new subject decoding. We also find that the pretrained PopT weights themselves reveal interpretable patterns for neuroscientific study. We show how the pretrained weights can be probed for connectivity and how the fine-tuned attention weights can be used to map task-specific functional salience.
@@
    a generic self-supervised learning framework, Population Transformer, PopT, that learns joint representations of varying spatially-sparse time series on top of pretrained single-time-series representations,
@@
    a demonstration that self-supervised pretraining systematically yields better performance, sample efficiency, and compute efficiency when downstream decoding aggregations of electrode embeddings, even for subjects held out during pretraining,
@@
    a trained usable off-the-shelf model that computes population-level representations of intracranial neural recordings.
@@ 2 Related Work
Self-supervised learning on neural data Channel independent pretrained models are a popular approach for neural spiking data [11], intracranial brain data [8, 12], and general time-series [13]. Additionally, in fixed-channel neural datasets, approaches exist for EEG [14, 15, 16], fMRI [17, 18, 19], and calcium imaging [20] datasets. However, all of this work do not learn population-level interactions across datasets with different recording layouts due to the single-channel focus or the ability to assume fixed-channel setups. Several works pretrain spatial and temporal dimensions across datasets with variable inputs [21, 22, 23, 24, 25], but most simultaneously learn the temporal embeddings with the spatial modeling, which make them challenging to interpret and computationally expensive to train. As far as we know, we are the first to study the problem of building pretrained channel aggregation models on top of pre-existing temporal embeddings trained across datasets with variable sampling of input channels, allowing for modeling of high quality (>2kHz sampling rate) intracranial data.
@@
Modeling across variable input channels Modeling spatial representations on top of temporal embeddings have been found to be beneficial for decoding [3, 7, 26], but prior works use supervised labels, so do not leverage large amounts of unannotated data. The brain-computer-interface field has been studying how to align latent spaces [27, 28, 29, 30, 31] which either still requires creating an alignment matrix to learn across datasets or only provides post-training alignment mechanisms rather than learning across datasets. Other approaches impute missing channels or learn latent spaces robust to missing channels [32, 33, 34], but these are more suited for the occasional missing channel rather than largely varying sensor layouts. We directly learn spatial-level representations using self-supervised learning across datasets to leverage massive amounts of unannotated intracranial data.
@@
Figure 1: Schematic of our approach. The inputs to our model (a) are the combined neural activities from a collection of intracranial electrodes in a given time interval. These are passed to a frozen temporal embedding model, which produces a set of time-contextual embedding vectors (yellow). The 3D position of each electrode (red) is added to these vectors to produce the model inputs (orange). The PopT produces space-contextual embeddings for each electrode and a [CLS] token (blue), which can be fine-tuned for downstream tasks. During pretraining, (b) the PopT is trained on two objectives simultaneously. In the first, the PopT determines whether two different sets of electrodes (orange vs brown) represent consecutive or non-consecutive times. In the second objective, the PopT must determine whether an input channel has been replaced with activity at a random other time that is inconsistent with the majority of inputs.
@@ 3 Population Transformer Approach
In order to learn a subject-generic model of intracranial activity that can handle arbitrary configurations of electrodes, we design a self-supervised training scheme that requires the model to learn representations of individual electrodes as well as groups of electrodes. One component of our self-supervised loss requires the model to identify which channels have been swapped with activity from the same channel, but at a different time point. To do this task, the model must build a representation of the channel’s activity that is sensitive to the context of all the surrounding channels. The other component requires the model to discriminate between randomly selected subsets of electrodes to determine if their activity has occurred consecutively in time or not. This requires the same sensitivity to context, but at the ensemble level. One can think of this swap and discriminate objective as exposing the model to many in-silico ablations of the brain, and asking the model to learn the connections between regions in the presence of these ablations.
A key aspect of our method is the fact that our objective is discriminative, rather than reconstructive, as is often the case in self-supervision [35, 8]. We found this to be necessary, because in practice, the temporal embeddings often have low effective dimension (see [8]), and reconstruction rewards the model for overfitting to “filler” dimensions in the feature vector (see Section 5).
We take additional steps to make our model subject and configuration generic. We provide the absolute position of every electrode to the model, which allows the model to learn a common position embedding space across subjects. We also vary the size of the subsets we select in our sampling procedure to ensure that the model can handle ensembles of differing number, which is important for neuroscience applications, in which experiments have varying number of electrodes, and analysis may be done on the electrode, wire, region, or brain level. Finally, we select that subsets are disjoint, to ensure that the model does not learn to solve the task by trivial copying.
@@
Architecture A schematic of our Population Transformer (PopT) approach is shown in Figure 1. Consider a given subject with $N_{c}$ channels indexed by $C=\{1,...,N_{c}\}$. Activity from channel $i$ at time $t$ can be denoted by $x^{t}_{i}$. The PopT takes as input an interval of brain activity $X=\{x^{t}_{i}|i\in C\}$ from a given time $t$ and a special [CLS] token. Per channel, each interval of brain activity is passed through a temporal embedding model $T$, in our case BrainBERT, to obtain a representation of each channel’s temporal context.
@@
Before being inputted to the PopT, each channel’s 3D position is added to this embedding, so the final input is $X_{B}=\{T(x)+pos(i)+\mathcal{N}(0,\sigma)|x\in X\}$. Here, we add Gaussian fuzzing to prevent overfitting to a particular set of coordinates. Spatial location is given by the electrode’s Left, Posterior, and Inferior coordinates [36]; see [8] for details on how these were obtained. Each coordinate is encoded using sinusoidal position encoding [9]. And the three encodings are concatenated together to form the position embedding $pos(i)=[e_{\text{left}};e_{\text{post.}};e_{\text{inf}}]$.
@@
The core of PopT consists of a transformer encoder stack (see Appendix A: Architectures). The output of the PopT are spatial-contextual embeddings of the channels $Y=\{y_{i}\}$ as well as an embedding of the CLS token $y_{cls}$. During pretraining, the PopulationTransformer additionally is equipped with a linear layer head for the [CLS] token output and separate linear layer heads for all other individual token outputs. These produce the scalars $\tilde{y}_{cls}$ and $\tilde{y}_{i}$ and respectively, which are used in the objective (see Figure 1a).
Pitch Volume Sent. Onset Speech/Non-speech
BrainBERT: single channel $0.53\pm 0.05$ $0.58\pm 0.08$ $0.68\pm 0.04$ $0.66\pm 0.09$
Linear + BrainBERT $0.59\pm 0.08$ $0.66\pm 0.08$ $0.70\pm 0.09$ $0.71\pm 0.11$
Deep NN + BrainBERT $0.58\pm 0.08$ $0.67\pm 0.08$ $0.71\pm 0.10$ $0.72\pm 0.10$

### v2

# Population Transformer: Learning Population-level Representations of Neural Activity

@@ Abstract
We present a self-supervised framework that learns population-level codes for arbitrary ensembles of neural recordings at scal. We address two key challenges in scaling models with neural time-series data: sparse and variable electrode distribution across subjects and datasets. The Population Transformer (PopT) stacks on top of pretrained representations and enhances downstream decoding by enabling learned aggregation of multiple spatially-sparse data channels. The pretrained PopT lowers the amount of data required for downstream decoding experiments, while increasing accuracy, even on held-out subjects and tasks. Compared to end-to-end methods, this approach is computationally lightweight and more interpretable, while still retaining competitive performance. We further show how our framework is generalizable to multiple time-series embeddings and neural data modalities. Beyond decoding, we interpret the pretrained PopT and fine-tuned models to show how they can be used to extract neuroscience insights from massive amounts of data. We release our code as well as a pretrained PopT to enable off-the-shelf improvements in multi-channel intracranial data decoding and interpretability.
@@ 1 Introduction
Building effective representations of neural data is an important tool in enabling neuroscience research. We are particularly interested in modeling neural activity recorded from spatially varying electrode layouts, such as intracranial recordings, which use invasive probes placed within the brain to provide high temporal resolution recordings of local neural activity (Parvizi & Kastner, 2018; Herff et al., 2020). But because of dispersed electrode placement within the brain volume, intracranial recordings (iEEG) suffer from data sparsity, and there is significant variability in probe placement across subjects (Parvizi & Kastner, 2018; Herff et al., 2020), leading to high variability in the semantics of the input channels. While these difficulties are particularly relevant for iEEG, they also exist for other neural data modalities such as non-invasive EEG due to variation in electrode layouts. Furthermore, annotated training data is often scarce, since it is limited by patient availability and labor-intensive labeling. This is a problem for many existing approaches which rely on supervised learning (Faezi et al., 2021; Herff et al., 2020; Martin et al., 2018; Metzger et al., 2023; Willett et al., 2023).
@@
To improve decoding data-efficiency, self-supervised pretraining on unannotated data can be employed to first learn generic representations of the recordings, independent of the downstream decoding task that requires task labels. This means that the model does not have to use valuable annotated samples to learn how to do feature extraction before it can do classification, improving the reach of neuroscientific research. In this paper, we are interested in developing generic representations of multi-channel intracranial recordings that enable efficient adaptation to a wide range of downstream decoding tasks. Prior work has shown how to pretrain subject-specific (Le & Shlizerman, 2022) or channel-specific (Wang et al., 2022) models of intracranial data, but such techniques ignore inter-channel relationships or commonalities that might exist across subjects. The most general approach would be to pretrain using data from multiple datasets, but would require tackling the aforementioned challenges of sparse electrode coverage and variable electrode placement between subjects.
@@
We propose Population Transformer (PopT), a self-supervised pretraining approach that learns subject-generic representations of arbitrary electrode ensembles. Transformers offer the flexibility to learn aggregate information across channel configurations, but massive data is needed to train the attention weights (Devlin et al., 2019). During pretraining, we train on large amounts of unannotated data and simultaneously optimize both a channel-level and ensemble-level objective. This requires the model to (1) build subject-generic representations of channel ensembles and (2) meaningfully distinguish temporal relationships between different ensembles of channels.
@@
Our PopT approach is modular, and builds on top of powerful single-channel temporal embeddings, which provides two key advantages. First, by separating the single-channel embedding and multi-channel-aggregation into different modules, we make our approach agnostic to the specific type of temporal embedding used, leaving room for future independent improvements along either the temporal or spatial dimension (an approach that has been validated in video modeling (Arnab et al., 2021)). Second, by taking advantage of learned channel embeddings, PopT training is computationally lightweight compared to their end-to-end counterparts (Appendix B) and baseline aggregation approaches (Figure 4), allowing for adoption in lower compute resource environments.
@@
Empirically, we find that our pretrained PopT outperforms commonly used aggregation approaches (Ghosal & Abbasi-Asl, 2021), and is competitive with end-to-end trained methods (Zhang et al., 2024; Yang et al., 2024; You et al., 2019). Moreover, we find that these benefits hold even for subjects not seen during pretraining, indicating its usefulness for new subject decoding. We also show that the pretrained PopT weights themselves reveal interpretable patterns for neuroscientific study. Finally, we demonstrate that our proposed framework is agnostic to the underlying temporal encoder further allowing it to adapt to other neural recording modalities.
@@
    a generic self-supervised learning framework, Population Transformer (PopT), that learns joint representations of arbitrary channel ensembles across neural datasets,
@@
    a demonstration that pretraining systematically improves ensemble representations for downstream decoding even for held-out subjects,
@@
    a trained and usable off-the-shelf model that computes population-level representations of high temporal resolution intracranial neural recordings.
@@ 2 Related Work
Self-supervised learning on neural data Channel independent pretrained models are a popular approach for neural spiking data (Liu et al., 2022), intracranial brain data (Wang et al., 2022; Talukder & Gkioxari, 2023), and general time-series (Talukder et al., 2024). Additionally, in fixed-channel neural datasets, approaches exist for EEG (Chien et al., 2022; Kostas et al., 2021; Yi et al., 2023), fMRI (Thomas et al., 2022; Kan et al., 2022; Ortega Caro et al., 2023), and calcium imaging (Antoniades et al., 2023) datasets. However, these approaches do not learn population-level interactions across datasets with different recording layouts, either due to a single-channel focus or the assumption that the channel layout is fixed. Several works pretrain spatial and temporal dimensions across datasets with variable inputs (Zhang et al., 2024; Yang et al., 2024; Jiang et al., 2024; Ye et al., 2024; Cai et al., 2023), but most simultaneously learn the temporal embeddings with the spatial modeling, which make them challenging to interpret and computationally expensive to train, especially for high temporal resolution signals. To our knowledge, we are the first to study the problem of building pretrained channel aggregation models on top of pre-existing temporal embeddings trained across neural datasets with variable channel layouts, allowing for modeling of high quality neural data.
@@
Modeling across variable input channels Modeling spatial representations on top of temporal embeddings has been found to be beneficial for decoding (Faezi et al., 2021; Le & Shlizerman, 2022; Azabou et al., 2024), but prior works use supervised labels, so do not leverage large amounts of unannotated data. The brain-computer-interface field has studied how to align latent spaces (Pandarinath et al., 2018; Karpowicz et al., 2022; Degenhart et al., 2020; Jude et al., ; Ma et al., 2023) which either still requires creating an alignment matrix to learn across datasets or only provides post-training alignment mechanisms rather than learning across datasets. Other approaches impute missing channels or learn latent spaces robust to missing channels (Talukder et al., 2022; Zhang et al., 2021; Chau et al., 2024), but these are more suited for the occasional missing channel rather than largely varying sensor layouts. We directly learn spatial-level representations using self-supervised learning across datasets to leverage massive amounts of unannotated intracranial data.
@@
Figure 1: Schematic of our approach. The inputs to our model (a) are the combined neural activities from a collection of intracranial electrodes in a given time interval. These are passed to a frozen temporal embedding model (dotted red outline: BrainBERT shown), which produces a set of time-contextual embedding vectors (yellow). The 3D position of each electrode (red) is added to these vectors to produce the model inputs (orange). The PopT produces space-contextual embeddings for each electrode and a [CLS] token (blue), which can be fine-tuned for downstream tasks. During pretraining, the PopT is trained on two objectives simultaneously. In the first, (b) the PopT determines whether two different sets of electrodes (orange vs brown) represent consecutive or non-consecutive times. In the second objective, (c) the PopT must determine whether an input channel has been replaced with activity at a random other time that is inconsistent with the majority of inputs.
@@ 3 Population Transformer Approach
Figure 1 overviews our Population Transformer (PopT) approach. The key ideas are: (1) to learn a generic representation of neural recordings that can handle arbitrary electrode configurations; and (2) to employ a modular system design that uses a transformer architecture to aggregate information from existing per-channel temporal embeddings. To do so, we employ a self-supervised pretraining approach to learn ensemble and channel level representations. Afterwards, one can fine-tune PopT on downstream decoding tasks. In addition to offering strong decoding results, including generalization to new subjects with different electrode configurations than training subjects (see Section 5), the modular system design is computationally lightweight (see Appendix B), can benefit from improved temporal representations, and is more readily interpretable (see Section 6).
@@
Architecture A schematic of our Population Transformer (PopT) approach is shown in Figure 1. We adapt a transformer backbone due to its ability to accommodate variable channel configurations. Consider a given subject with $N$ channels indexed by $C=\{1,...,N_{c}\}$, and an arbitrary subset of channels $S\subseteq C$. Activity from channel $i$ at time $t$ can be denoted by $x^{t}_{i}$. The PopT takes as input an interval of brain activity $X^{t}=\{x^{t}_{i}|i\in S\}$ from a given time $t$ and a special [CLS] token. Per channel, each interval of brain activity is passed through a temporal embedding model $B$, in the figure’s case BrainBERT, to obtain a representation of each channel’s temporal context.
@@
To allow the model to learn a common brain state representation across layouts, each channel’s embedding is augmented with its 3D position, so that the final input to the PopT is $X^{t}_{B}=\{B(x^{t}_{i})+pos(i)+\mathcal{N}(0,\sigma)|x^{t}_{i}\in X^{t}\}$. Spatial location is given by the electrode’s Left, Posterior, and Inferior coordinates for iEEG electrodes (Wideman, 2024), and XYZ positions for EEG electrodes. We add Gaussian fuzzing to each coordinate location to prevent overfitting to a particular set of coordinates. Each coordinate is then encoded using sinusoidal position encoding (Vaswani et al., 2017). Membership in a particular ensemble (see below: ensemble-wise loss) is also encoded. The four encodings are concatenated together to form the position embedding $pos(i)=[e_{\text{left}};e_{\text{post.}};e_{\text{inf}};e_{\text{ensemble}}]$.
@@
The core of PopT consists of a transformer encoder stack (see Appendix A: Architectures). The output of the PopT are spatial-contextual embeddings of the channels $Y=\{y_{i}\}$ as well as an embedding of the CLS token $y_{cls}$. During pretraining, the PopulationTransformer additionally is equipped with a linear layer head for the [CLS] token output and separate linear layer heads for all other individual token outputs. These produce the scalars $\tilde{y}_{cls}$ and $\tilde{y}_{i}$ respectively, which are used in the pretraining objective (see Figure 1b and c).
Self-supervised loss Our loss function has two discriminative components: (1) ensemble-wise — the model determines if activities from two channel ensembles occurred consecutively, requiring an effective brain state representation at the ensemble-level, (2) channel-wise — the model identifies outlier channels that have been swapped with a different timepoint’s activity, requiring sensitivity to surrounding channel context.
A key aspect of our method is the fact that our objective is discriminative, rather than reconstructive, as is often the case in self-supervision (Liu et al., 2021; Wang et al., 2022). We found this to be necessary, because in practice, the temporal embeddings often have low effective dimension (see Wang et al. (2022)), and reconstruction rewards the model for overfitting to “filler” dimensions in the feature vector (see Section 5).
Pretraining In ensemble-wise discrimination (fig. 1b), two different subsets of channels $S_{A},S_{B}\subset C$ are chosen with the condition that they be disjoint $S_{A}\cap S_{B}=\emptyset$. During pretraining, the model receives the activities from these channels at separate times $X_{A}^{t}=\{x^{t}_{i}\mid i\in S_{A}\}$ and $X_{B}^{t^{\prime}}=\{x^{t^{\prime}}_{i}\mid i\in S_{B}\}$. The objective of the task is then to determine whether these states $X_{A}^{t}$ and $X_{B}^{t^{\prime}}$ have occurred consecutively in time or are separated by some further, randomly selected interval. Given the output of the classification head, the loss function $\mathcal{L}_{N}$ is again the binary cross entropy. We also vary the number of input channels during sampling to ensure the model handles ensembles of different sizes.
Additionally, we select disjoint subsets for ensemble-wise discrimination to prevent the model from solving tasks through trivial copying.
In channel-wise discrimination (fig. 1c), the model must determine whether a channel’s activity has been swapped with activity from a random time. Precisely, activity from each channel $i$ is drawn from a time $t_{i}$. All channels are drawn from the same time $t_{i}=T$, and then 10% of the channels are randomly selected to have their activity replaced with activity from the same channel, but taken from a random point in time $t_{i}\neq T$. Then, given the token outputs of PopT, the channel-wise loss function $\mathcal{L}_{C}$ is the binary cross entropy. Then, our complete objective function is $\mathcal{L}=\mathcal{L}_{N}+\mathcal{L}_{C}$.
Fine-tuning During fine-tuning, the [CLS] intermediate representation, $\tilde{y}_{cls}$ of the pretrained PopT is passed through a single layer linear neural network to produce a scalar $\hat{y}_{cls}$; this forms the input to the binary cross entropy loss for our decoding tasks (see Section 4).
4 Experiment Setup
Data We use two types of neural time-series data: intracranial and scalp electroencepholography (iEEG and EEG). Whereas iEEG probes are surgically implanted within the 3D brain volume, EEG electrodes lie on the scalp. These cover two resolution extremes of neural time-series data modalities.
iEEG: We use the publicly available subject data from Wang et al. (2022). Data was collected from 10 subjects (total 1,688 electrodes, with a mean of 167 electrodes per subject) who watched 26 movies while intracranial probes recorded their brain activity. Of the sessions, 19 are used for pretraining, and 7 of the sessions are held-out for evaluation.
EEG: We use the Temple University Hospital EEG and Abnormal datasets (TUEG and TUAB (Obeid & Picone, 2016)) for pretraining and task data respectively. We remove all task subjects from the pretraining set and follow the data preprocessing practices in Yang et al. (2024); Jiang et al. (2024).
Decoding Tasks We evaluate on 5 different classification tasks: 4 auditory-linguistic tasks used in the evaluation of Wang et al. (2022) and 1 widely evaluated seizure detection task from Obeid & Picone (2016). Of the auditory-linguistic tasks, two of the tasks are audio focused: determining whether a word is spoken with a high or low pitch and determining whether a word is spoken loudly or softly. And two of the tasks have a more linguistic focus: determining whether the beginning of a sentence is occurring or determining whether any speech at all is occurring. The EEG TUAB seizure detection task is a binary classification of pathological or normal EEG recording.
To test scaling with arbitrary ensemble sizes, we select subsets of electrodes based on their individual linear decodability, with the smallest subsets containing the electrodes with highest decodability.
Baselines For controlled baselines, we concatenate the single-channel temporal embeddings and train a linear (Linear) or non-linear (Deep NN) aggregator on the decoding task. These enable us to directly assess how much PopT improves upon existing aggregation approaches (Ghosal & Abbasi-Asl, 2021). These approaches cannot be pretrained across subjects due to the changing meaning and quantity of inputs. To test the effectiveness of pretraining, we also compare against a non-pretrained PopT. Methods compared For the iEEG experiments, we also compare against Brant (Zhang et al., 2024), which is an end-to-end iEEG trained encoder. We take the fully pretrained Brant model, and fine-tune on our iEEG tasks combining channels with linear aggregation. For the EEG experiments, we compare against BIOT (Yang et al., 2024) and LaBraM (Jiang et al., 2024). Both BIOT and LaBraM use discrete channel ID as positional encodings, which only allows them to be used for layouts with discete channel locations such as EEG. They also train end-to-end in contrast to our modular approach.
Model Pitch Volume Sent. Onset Speech/Non-speech
BrainBERT:
      Linear Agg. $0.59\pm 0.08$ $0.66\pm 0.08$ $0.70\pm 0.09$ $0.71\pm 0.11$
      Deep NN Agg. $0.58\pm 0.08$ $0.67\pm 0.08$ $0.71\pm 0.10$ $0.72\pm 0.10$